AlbIn for MUI

David von Plenker-Tind Sirius Design March 26, 1995

Contents

1 Introduction 3 1.1 What is this? . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Sirius Design . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.5 \begin{matrix}c \\ ◯ \end{matrix} Copyrights . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Sirius Design 7 2.1 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3 Tutorial 9

3.1 Quick-start . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.1 Loading the first database . . . . . . . . . . . . . . . 9

4 The Windows 11 4.1 General information . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Album Window . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3 Song Window . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.4 Album Comments Window . . . . . . . . . . . . . . . . . . 13 4.5 Song Comments Window . . . . . . . . . . . . . . . . . . . 14 4.6 Artist Info Window . . . . . . . . . . . . . . . . . . . . . . . 15 4.7 List Window . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.8 Find Window . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.9 Export Window . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.10 Prefs Window . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5 The Menues 19 5.1 Project Menu . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.2 Sort Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.3 Album Menu . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.4 Misc Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.5 Open/Activate Menu . . . . . . . . . . . . . . . . . . . . . . 21 5.6 Artist-Title List Menu . . . . . . . . . . . . . . . . . . . . . 22

6 Various information 23 6.1 Erland Rating . . . . . . . . . . . . . . . . . . . . . . . . . . 23 6.2 MUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

Chapter 1

Introduction

1.1 What is this?

This is a manual for AlbIn MUI in TEX-format. This dvi-file gives you a nice output from any plattform. This manual should explain everything you have to know about AlbIn MUI, and how you use it in the most efficient way.

Read the Quick-start to get going with AlbIn MUI right away.

1.2 Sirius Design

Sirius Design consists of three programmers. We are making high-quality software for the Amiga computer. These programs are easy to use and have a vast range of functions when needed.

Sirius Design consists of

Niklas Dahlman Mathias Svensson David von Plenker-Tind
Ympningsvägen 14 Ympningsvägen 10 Priorvägen 14
S-240 14 Veberöd S-240 14 Veberöd  S-240 10 Dalby
d92nd@efd.lth.se  ms@p15.dennis.ct.se dat94pvo@ludat.lth.se

1.3 Requirements

You have the following software and hardware to use AlbIn MUI:

OS KickStart (ROM Libraries) v37+

Libs: reqtools.library v38+ muimaster.library v2.2+

Also: the tool sys:tool/printfiles a t: assign

1.4 Installation

With AlbIn MUI a installer-script is provided. This script uses the tool ’installer’ from Commodore. If you have Wb2.1+ you will probably already have this tool, otherwise it’s on the fish-series, the Aminet or in a good PDlibrary.

It’s easy to install AlbIn MUI with it’s install-script. If you don’t have the installer the easiest way to make the installation is to make it manually. Just copy AlbIn MUI and AlbIn MUI.guide and there icons to the directory where you want it. If you want to have the example databases installed you only have to copy the intire directory to the same directory as the executable.

1.5 \begin{matrix}c \\ ◯ \end{matrix} Copyrights

Under this section, all the copyright notices will be together with the disclaimer that you must have in your manual.

AmigaGuide, AmigaGuide.info, amigaguide.library (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved.

Installer and Installer project icon (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved.

reqtools (c) Copyright 1991-94 Nico François, Inc. All Rights Reserved. THIS SOFTWARE IS PROVIDED ”AS IS”. NO REPRESENTA-

TIONS OR WARRANTIES ARE MADE WITH RESPECT TO THE AC-CURACY, RELIABILITY, PERFORMANCE, CURRENTNESS, OR OP-ERATION OF THIS SOFTWARE, AND ALL USE IS AT YOUR OWN RISK. NEITHER COMMODORE NOR THE AUTHORS ASSUME ANY RESPONSIBILITY OR LIABILITY WHATSOEVER WITH RESPECT TO YOUR USE OF THIS SOFTWARE.

1.5 \begin{matrix}c \\ ◯ \end{matrix} Copyrights  5

AmigaGuide, AmigaGuide.info, amigaguide.library, WDisplay, WDisplay.info (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore.

AMIGAGUIDE SOFTWARE IS PROVIDED ”AS-IS” AND SUBJECT TO CHANGE; NO WARRANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPONSIBILITY IS ASSUMED.

Installer and Installer project icon (c) Copyright 1991-93 Commodore-Amiga, Inc. All Rights Reserved. Reproduced and distributed under license from Commodore.

INSTALLER SOFTWARE IS PROVIDED ”AS-IS” AND SUBJECT TO CHANGE; NO WARRANTIES ARE MADE. ALL USE IS AT YOUR OWN RISK. NO LIABILITY OR RESPONSIBILITY IS ASSUMED.

The Installer project icon was redesigned

6  Introduction

Sirius Design

2.1 Distribution

AlbIn is freely distributable, but if you like it and use it for a longer period of time, then we would be very happy if you respect the shareware fee and send us some money.

AlbIn may be distributed as an archived to save space. Make sure that all of the files below are included in the arcive. Please make no changes in the code, catalogs, documents or exampel files without contacting us first!

When you give AlbIn to your friends make sure you include:

AlbIn MUI/AlbIn MUI Databases/Andras Skivor.AlbIn MUI AlbIn MUI/AlbIn MUI Databases/Andras Skivor.AlbIn MUI.info AlbIn MUI/AlbIn MUI Databases/Davids Skivor.AlbIn MUI AlbIn MUI/AlbIn MUI Databases/Davids Skivor.AlbIn MUI.info AlbIn MUI/AlbIn MUI Databases/Depeche.AlbIn MUI AlbIn MUI/AlbIn MUI Databases/Depeche.AlbIn MUI.info AlbIn MUI/AlbIn MUI Databases/Niklas Skivor.AlbIn MUI AlbIn MUI/AlbIn MUI Databases/Niklas Skivor.AlbIn MUI.info

AlbIn MUI/AlbIn MUI Exports/Niklas Skivor.guide AlbIn MUI/AlbIn MUI Exports/Niklas Skivor.guide.info

AlbIn MUI/AlbIn MUI AlbIn MUI/AlbIn MUI.info AlbIn MUI/AlbIn MUI.guide AlbIn MUI/AlbIn MUI.guide.info

8  Sirius Design

Tutorial

3.1 Quick-start

Ok. Let’s get started right away. To start AlbIn MUI you only have to doubble-klick on the icon. Hmmm, there’s just one thing, you have to have the very nice software-pack ’MUI’ by Stefan Stuntz installed on your system (see p. 24).

When your system is all set up (the program installed) it’s time to get started. Doubble-click on the nice icon (Plain or MagicWB doesn’t matter, they both launch the program) and wait for a while if this is the first MUI application since the last reboot. If you haven’t configured MUI to open on a different screen than Workbench, a window should appear on the WB-screen. You can have AlbIn MUI on a public screen, but then you will loose the application window feature (so you can drag-n-drop your databases on the AlbIn MUI) windows.

3.1.1 Loading the first database

To get acuinted with the windows I suggest that you load an already saved database. Select the <example.AlbIn MUI> and choose Ok in the file requester to load the file. Now you will see the slider at the bottom of the window to shrink and the number to the right will increase. It may take a while if you have a slower machine (If you think it’s to slow it’s only to by an accelerator card or a faster computer!!). I think it’s rather fast on my A3000 with 8 colors in hires interlace and it’s wizing fast on a standard 4 color hires screen. If you run AlbIn MUI on a hires screen you might notice that the windows take up a lot of space. Well, the only way to ”shrink” the windows is to run it in hires interlace or another display mode with the same or higher resolution. Yes, I know it flickers like hell, but wear a pair of dark sunglasses or buy a flicker fixer and a real monitor… (the sunglasses are probably cheaper but the other solution will be better in the long run).

The Windows

4.1 General information

The operation of the GUI complies to the MUI standard. This means that all standard shortcuts like >TAB< for changing gadget and >Alt< + >TAB< for stepping through windows. The order of the tab cycle is defined to offer the most efficient way to enter data into the database. The Song Window is a good example. When you enter a title of a song and press >ENTER< the ’Add Time’ gadget gets activated. Pressing >ENTER< again and the requester pops up and asks you to enter the time of the song, and when you press >ENTER< tiwce (one to accept the input in the string gadget and one to accept the requester) the ’New Song’ gadget is activated again, making it easy to enter the next song without fiddling with alt+tab or grasping for the mouse.

4.2 Album Window

This is the window that opens first. The window contains some important strings and a ”format”-cycle gadget. It also contains several orientation gadgets that you use to move around in your register.

Artist Enter the name af the artist/group here. When you enter something here the titles in the other windows will change automatically. The ask-sign at the end of the string gadget is connected to the ’Artist List’ window (see section 4.6). The window opens and show the groupmembers if you have entered any together with another album.

Title Enter the title of the LP/CD/MC etc. The titles on the other windows will change here too.

PlayTime The total time of the album.

Edit/View Comment... Brings up a window with more string gadgets. See also 4.3.

Edit/View Songs... Brings up a window with info about the tracks on the album. See also 4.5.

New Makes a new post at the end of the register.

Insert Inserts a new post in the register. The post will loose its position when you sort the albums.

Delete... Deletes the selected post from the register. A requester will show up and warn you. You can’t get back what you have deleted.

Slider With the slider you can directly select the position in your database that you like. You can also use the arrow images on the right side of the slider to move one step up/down.

Format This cycle gadget shows the format (media) of your post. The following medias are represented: CD – Compact Disc CDS – CD Single CDM – CD Maxi MCD – Maxi CD CD+G – CD + Graphics MD – Mini Disc LD – Laser Disc LP – Long Play Single Maxi EP – Extended Play MC – Music Cassette DAT – Digital Audio Tape DCC – Digital Compact Cassette CD-ROM – Compact Disc - Read Only Memory Video Other

4.3 Song Window

The Song Window consists of a rather large listview gadget in which the different song on the album is shown. The time of the songs are also displayed in the listview gadget. On the right side of the window there are some control gadgets to make the life some what simpler.

Listview gadget This is where you enter the tracks of the album.

Add Song If you press ’Add Song’ a new entry will be made in the listview gadget. Enter the title of the track. A total of 50 characters (time included).

Add Time This button brings up a requester asking for the time of the selected track. Enter the time with your own separators.

Delete Deletes the selected track from the list. A requester will notify you before you delete a track.

Up Moves the selected track one position up in the list.

Down Moves the selected track one position down.

Edit/View Comment... Opens the ’Album Comments’ window.

GetFile Gadget Select the lyrics for the album. The lyrics should be a normal ascii file, but it can also be a .guide-file if you use Multiview or Amigaguide as a viewer. How you wish to type it in, or how other has typed it in, is of your own choice. You select your viewer in the prefs window 4.10.

Edit Lyrics Press this button and you will get your own editor (previous selected in the prefs window 4.10). Enter your lyrics and save the file as plain ascii. Choose the file with the getfile gadget and view with the…

View Lyrics … button. Select the viewer in the icon and the lyrics will appear after pressing this button.

4.4 Album Comments Window

This window shows the different comments about your album. You enter the information in the string gadgets and press >TAB< to proceed to the next string gadget. MUI makes it possible to go to the next string gadget by pressing >ENTER< but that is not what the style guide says…

The following information can be stored:

Record Company Which company who has published the album. Made in Which country in which the album has been made in. Cat. Nr. Short for ’Catalog Number’ – the number from the record company

Cover Design The name(s) of the people(s) responsible for the album cover. Release Date When the album was made available for the public.

Speed Rotation speed of the record (only old vinyls).

Recording Technology Applying to digital media such as the CD, the Mini Disc, the DAT

and the DCC.

Video Format Which type of tape the video was recorded on.

Rating If the record has been reviewed in a magazine you can type in the rating here. You can also make your own rating scale and type it in here. See p. 23 for more information about the ’Erland rating’ system.

Tape Quality Dolby type and cassette quality for

Cover If you have scanned your covers you can choose a path to the directory where you keep the pictures and you can use your favourite viewer (like Viewtek) to show the pictures. You select the viewer in the preferences (see 4.10

4.5 Song Comments Window

This is the window were you enter your comments about the lyrics, the music and the post production. There are a lot of string gadgets here to, so I write a breif comment about each of them:

Music The writer of the music. Lyrics The writer of the lyrics. Vocals Vocalists – can also be arranger.

Mixed by The person who mixed the album. Produced by The producer of the album. Engineered by The recording engineer.

Mixed in The studio or other facility in which the album were mixed.

Recorded in The studio where the album were recorded at.

Published by Publisher of the album – if other than the record company. There could be different publishers in different countries.

Other Other comments about the album. References to other groups or albums etc.

4.6 Artist Info Window

This windows consists mainly of a listview gadget that contains the group members.

4.7 List Window

All albums are listed here together with thier format. This list reflects the sorting (or lack of such).

4.8 Find Window

I you are lycky you might have so many albums that searching for something special might be tiresome if you do it manually. Well, you can always use the find function in AlbIn MUI to find what you search for. When you open the window, you will see a lot of checkmark gadgets. If you tick them, AlbIn MUI will search that field in every album. The checkmarks represent every string gadget were you can enter text. There are also a short cut to mark all items wich may contain a name of a person or a title of a song or an album. There’s also a gadget to clear all checked items when you want to make a different search. You can also ignore the casing of the words that you search for which means that if you search for ’Energy Rekords’, ’energy rekords’ as well as ’eNerGY rEKorDs’ will be found. If you haven’t marked the ’Ignore Case’ gadget AlbIn MUI will not find anything.

You can export the search result as an AlbIn MUI database. A file requester will pop up so you can enter the name of the new archive.

4.9 Export Window

This is where you export your database in different formats. You can export to three ’destinations’. They are:

1. Preferences printer

2. ASCII

3. Guide

If want to use the first option, you have to check your preferences settings to ensure that you use the right printer driver. AlbIn MUI will perform a form-feed between every entry in your database if have chosen the option to export an entire entry so it’s easier to seperate the entries. If you only export a list over titles or artists, then AlbIn MUI will print everything in one piece. The ’prefs printer’ option is the same as the ASCIIexport option, except you get your result on a paper instead of a file.

If you export ASCII then you can set the default export directory in the preferences window 4.10. This also applies to the guide-export option.

4.10 Prefs Window

This is where you enter the preferences that has to do with AlbIn MUI (the GUI preferences is set by the ’MUI’ prefs program). The default viewers, editor and paths are set to be most standard as possible. The gadgets at the right are present so you can select almost everything with the mouse (you have to enter the options by hand).

Text viewer The command line for your favourite textviewer. The default is ’More’, hopefully present in the ’C’directory. Most text viewers can use the filename as an option so it can be read direct (without a file requester). The file name should be replaced by ’commandline and AlbIn MUI will insert the name of the file that you have specified. You may have to quote the name in case of spaces in the file name.

Picture viewer This is the command line for your picture viewer. The same thing applies to the viewer – ’used by AlbIn MUI to insert the filename. Default is ’Display’.

Text Editor This is the command line for the text editor. If you use CygnusEd or GoldEd you might have to use the option -sticky. I don’t know how Edge or TurboText is working so you have to consult your manual to see how you send file names as an option to the editor. Otherwise the same thing as above still applies to this command line. The standard editor is set to Ed (it’s not the top-of-the-line text editor but it’s free).

Text path The default path to your lyrics. If you don’t select a path it will be the current path.

Picture Path The default path to your cover pictures. Current path will be used if you don’t select a default path.

Text path The text path if you export ascii. The current path will be used if you don’t specify anything.

Guide path Export path for your giudes. Also set to current path if you don’t specify.

Load Dir The directory were you keep the archives to be loaded

Save Dir The directory were you save your archives (I don’t know why you would like different paths for saving and loading).

Save Options Two radio buttons for the save options. If you want to be asked to overwrite the old archive everytime you select the ’Save’ option you should choose the latter option, otherwise choose the first option and then you won’t be bothered by the anoying requester.

18  The Windows

The Menues

5.1 Project Menu

New Creates a new archive. If you have any unsaved changes AlbIn MUI will warn you. You can then either choose to clear the memory or save and then clear or cancel the action.

Open… AO A file requester is prompting you to select an archive to be loaded in to the memory. This version of AlbIn MUI is only using the RAM memory for storing the data. In later versions AlbIn MUI will be able to read/write the data directly off the drive where you store it.

Save AS Saves the archive with the current name. If you have choosen to ’Ask first’ in the ’prefs’ window then AlbIn MUI will always ask if you will overwrite the old archive with the same name, if you haven’t choosen that option, AlbIn MUI will simply overwrite the old archive without warning. If the archive doesn’t have a name, then this option will be the same as ’Save as’.

Save as… AA Let you save an archive with a different name. This option should also be used the first time (even if you don’t know about it). If an archive with the name that you selected then AlbIn MUI will ask you if you want to replace the old archive or give it another name (or you can cancel the action as always).

Export… AE Export the arcive to three different destinations. Read more about export under section 4.9.

About… A? Displays a requester with information about AlbIn MUI.

Quit… AQ Exit AlbIn MUI. If you have unsaved changes then AlbIn MUI will warn you before the exit. You can also cancel this action if you have these unsaved changes.

5.2 Sort Menu

When AlbIn MUI is adding an album (if you press ’New’ in the main window), it will be inserted after the last album in the current list of albums. To make the list sorted you will have to select sort method in the menu. An empty string will be sorted first.

Seperate Formats A0 If you want to sort every format of albums for them self, then this option should be marked. If you don’t mark this option the ’format’ cycle gadget will be ignored when sorting the albums.

Artist – Title A1 Sorting by the ’Artist’ field in the main window. The second criteria is the ’Title’ field in the same window.

Artist – Date A2 Sorting by the ’Artist’ field in the main window (just like above) and then by the ’Release Date’ field in the ’Album Comments’ window. It’s a good idea to have the year first and an additional date later so.

Title – Artist A3 Sorting with the ’Title’ first and then the ’Artist’.

Rec. Comp. – Date A4 Sorting by the ’Record Company’ field in the ’Album Comments’ window and then the ’Release Date’ field in the same window.

Rec. Comp. – Cat. Nr. A5 Sorting by the ’Record Company’ field in the ’Album Comments’ window and then the ’Catalouge Number’ in the same window.

Time – Artist A6 Sorting by the ’Play Time’ field in the main window and then by the ’Artist’ field in the same window.

5.3 Album Menu

New AN Makes a new album-entry in the list. This will be inserted last.

Insert AI Inserts a new album after the selected album.

Delete… AD Deletes the selected album. Arequester will ask you if you really want to delete this album.

Next Steps one album forward – to the next album in the list.

Prev Steps one album backward – to the previous album in the list.

First Jumps to the first item in the list over albums.

Last Jumps to the last item in the list over albums.

Artist List AR Opens the window with the list over albums. See section 4.7 for more detail.

5.4 Misc Menu

Guide Preview AP Generates a preview of the exported guide. The format is controlled by the ’Export’ window (see section 4.9).

Append Database Adds a previously saved database to the current database. N.B. Duplicates may occur. AlbIn MUI does not check for duplicates, so you have to remove them manually.

Find… Shows the ’Find’ window (see section 4.8. This window is also shown if you select ’Find Next’ without anything in the searchstring gadget.)

Find Next AF Finds the next item that matches the search criteria. It searches from the selected album to the end of the database.

Prefs… Shows the ’Preferences’ window (see section 4.10).

5.5 Open/Activate Menu

Album AB Opens the ’Album’ window. This is always open so this short cut is only activating the window. See also section 4.2 for further information about the window.

Song AG Opens the ’Song’ window. This is one of the two windws with a list gadget. See also section 4.3 for further information about the window.

Album Comments AU Opens the ’Album Comments’ window. See also section 4.4 for further information about the window.

Song Comments AC Opens the ’Song Comments’ window. See also section 4.5 for further information about the window.

Artist Info AT Opens the ’Artist Info’ window. This is the second window with a lista gadget. See also section 4.6 for further information about the window.

5.6 Artist-Title List Menu

The items in this menu is for the Artist-Title list only. You can select several records at one time by holding the left mouse-button down while draggging up or down.

Save Selected As… Makes it possible to save a part of the archive. Select some records and choose this option to save the records. You will be asked to give the new archive a name in the file-requester.

Move Selected Up A+ Moves the selected albums up. What’s it’s really doing is putting the record over the selection after instead.

Move Selected Down A- The same thing as above but reversed. Moves the selected albums down in the Artist-Title list.

Various information

6.1 Erland Rating

+5 Probably the all-time greatest album in the history of the universe. +4 Truly an all-time classic, one to buy and cherish forever. +3 A very good album indeed, and highly recommended. +2 A good album but with limited appeal or an album that has wider

appeal but is slightly less good. +1 A good album. Pleasant to listen to and certainly worth a listen.

0 Competent but unoutstanding. It’s not a *bad* record as such, it

just isn’t particularly *good* either -1 An album you should have reservations about. It’s not *really* bad,

in fact you’ll probably decide to keep it, but you probably wont play it that much.

-2 The sort of album which, unless you are completely fanatical about

the artist concerned you probably shouldn’t buy. -3 The sort of album which, even if you are completely fanatical about

the artist concerned you probably shouldn’t buy. -4 Really a very bad album indeed. Connosieurs of really bad albums

should buy this. -5 An album so bad you are unlikely ever to hear its like again.

6.2 MUI

MUI - MagicUser Interface Version 2.2 \begin{matrix}c \\ ◯ \end{matrix}Copyright 1993/94 by Stefan Stuntz

- ShareWare - MUI is an object oriented system to create and maintain graphical user

interfaces. From a programmers point of view, using MUI saves a lot of time and makes life much easier. Thinking about complicated terms like window resizing or font sensitivity is simply not neccesary.

On the other hand, users of MUI based applications have the ability to customize nearly every pixel of a programs interface according to their personal taste.

This distribution is interesting for both, users and programmers. Please have a look at the supplied demo programs and at the documentation to see what MUI has to offer.